OpenStudy (anonymous):

solve for x by using the exponential form: log3 (7-x)=3

OpenStudy (anonymous):

The first (and best) rule of logs is the definition \[log_b (a) = k \iff b^k = a\]

OpenStudy (anonymous):

In your case, what is a, b, and k?

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

not quite.

OpenStudy (anonymous):

I'll give you a hint. k = 3.

OpenStudy (anonymous):

oh the 7-x

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So what the definition means is if you have something that looks like this: \[log_b (a) = k\] It means the same thing as this: \[b^k = a\] And vice versa.

OpenStudy (anonymous):

So you have something that looks like this: \[log_3(7-x) = 3\] How would you rewrite it using the definition?

OpenStudy (anonymous):

3^3= 7-x

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

And solving from there is a breeze.

OpenStudy (anonymous):

yepp thanks!

OpenStudy (anonymous):

now a question in my textbook says:" use properities of logs or a definition to simpligy the expression: If f(x)= 10^x, find (log2). can you translate that for me?

OpenStudy (anonymous):

Huh? I'm not sure I understand what it's asking for.

OpenStudy (anonymous):

the log base 2 of 10^x ?

OpenStudy (anonymous):

i guess that is what it i asking?

OpenStudy (anonymous):

If f(x)= 10^x, find f(log2).

OpenStudy (anonymous):

Oh. that just means plug in log(2) for x. and since the base isn't specified it's log base 10.

OpenStudy (anonymous):

oh so it would be a decimal #?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[f(log(2)) = 10^{log(2)}\] Since \[log_{10} 2 = k \iff 10^k = 2\] Then \[10^{log(2)} = 10^k = 2\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

The take away message being if you raise the base to the power of a log you will get the input. \[b^{log_b(x)} = x\]

OpenStudy (anonymous):

Same for taking the log of a power.. \[log_b(b^x) = x\]

OpenStudy (anonymous):

But if you can't remember all those rules ( I rarely do ) you can always find them from the rule I gave at the start. The definition of the log.

OpenStudy (anonymous):

x^2= 10?

OpenStudy (anonymous):

No no. Lets say that \[log_{10}(2) = k\] What would that mean by the law of logs?

OpenStudy (anonymous):

Go back to the definition I gave at the beginning. What is a,b,and k?

OpenStudy (anonymous):

Actually in this case k is k, so no need for that ;p